Complete the ordered pairs for the equation 2x + y = 10.
(5, ), ( , 10), ( , –2), (0, )

Answer by ankor@dixie-net.com(22740)

Complete the ordered pairs for the equation 2x + y = 10.
(5, ), ( , 10), ( , –2), (0, )
:
Substitute the given value for x or y in the original equation, solve equation:
(5, __)
2x + y = 10
2(5) + y = 10
10 + y = 10
y = 10 - 10
y = 0
:
(__, 10)
2x + y = 10
2x + 10 = 10
2x = 10 - 10
x = 0/2
x = 0
:
(__, -2)
2x + y = 10
2x - 2 = 10
2x = 10 + 2
2x = 12
x = 12/2
x = 6
:
(0, __)
2(0) + y = 10
y = 10
:
Do you understand this now? Any questions?